OpenStudy (anonymous):

We are now in rationalizing and I do understand just a very bit of it. Help again, please. p^1/5 p^7/10 p^1/2 over (p^3)^-1/5 =?

OpenStudy (blockcolder):

\[\Large \frac{p^{1/5}p^{7/10}p^{1/2}}{(p^3)^{-1/5}}=\frac{p^{1/5+7/10+1/2}}{p^{-3/5}}=p^{7/5-(-3/5)}=p^2\]

OpenStudy (anonymous):

Thank you again :)

OpenStudy (blockcolder):

No problem. :D